In a packet of stickers there are 18 small stars, 5 big stars, 11 small rockets, and 7 big rockets. Kevin is going to choose one of these stickers from the packet at random to put on his artwork. What is the probability that the sticker Kevin chooses is big or is a rocket? Do not round intermediate computations, and round your answer to the nearest hundredth.
Found 2 solutions by ikleyn, ewatrrr:
Answer by ikleyn(52790)
Answer by ewatrrr(24785)


Hi
 41 stickers:   18ss, 5bs, 11sr, 7br
 P(b or r) = 12/41 + 18/41 - 7/41 = 23/41 = .56 (nearest hundredth)
Wish You the Best in your Studies.